AnnabelleHernandez:

Look at the following numbers: -10, -5, 0, 5 Which pair of numbers has a sum of 0? -10, 5 0, 5 5, -5 -5, 0

supie:

So basically you add up both of the numbers and see which one adds up to 0 So option 1 -10+5 = ? option 2 0+5=? option 3 5+-5=? option 4 -5+0=?

supie:

Any idea @AnnabelleHernandez?

AnnabelleHernandez:

yeah i think so

supie:

Ok so do you know the answer?

AnnabelleHernandez:

it's option 3

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AnnabelleHernandez it's option 3 \(\color{#0cbb34}{\text{End of Quote}}\) Yes

AnnabelleHernandez:

thx

supie:

np You can close this post now

AnnabelleHernandez:

ok

jhonyy9:

good job @supie

supie:

Thanks.